### v1

# Studying Workarounds in Software Forms: An Experimental Protocol

Abstract
Workarounds enable users to achieve goals despite system limitations but expose design flaws, reduce productivity, risk compromising data quality, and cause inconsistencies. We propose an experimental method to investigate how users employ workarounds when the data they want to enter does not align with software form constraints. By conducting user studies based on this method, we can analyze how workarounds originate and impact system design and data integrity. Understanding workarounds is essential for software designers to identify unmet user needs.
Keywords:
Workarounds Experimental Protocol User Study
1 Introduction
Software applications often require users to enter data for subsequent processing. Examples include a citizen applying for benefits, a banking assistant recording transaction details in a financial system, a sales representative reporting customer visits in a CRM platform, a physician documenting patient symptoms and diagnostics in a medical application, a customer filing an insurance claim, or a photographer inputting parameters to optimize image rendering in a photo editor. In most cases, these tasks are accomplished through forms, which serve as the primary interface for data entry.
In Information Systems, software applications often misalign with user needs and goals [15], leading them to improvise solutions. These improvised workarounds are deliberate deviations from intended software usage that emerge when application forms and workflows fail to accommodate user needs [2]. Sometimes, developers recognize workarounds as useful system features and integrate them in their designs.
The presence of workarounds can have positive and negative consequences [3]. Workarounds can reveal missing features, overlooked design flaws, or inefficient workflows. For example, users might develop complex external processes to generate a report, unaware of an existing built-in functionality for the task. While workarounds can be beneficial by enabling users to overcome system limitations, they also pose risks. Deliberate deviations may lead to unintended consequences, such as data loss, system inconsistencies, or compromised data quality. For instance, a user entering irrelevant data to bypass a required field might unintentionally affect the accuracy of downstream reports, resulting in semantic inconsistencies where the system operates technically correctly but the data becomes unreliable [27]. Understanding the implications of workarounds is thus critical for improving software design and usability.
We propose an experimental method for investigating how users develop workarounds when their goals do not align with software constraints. This method enables descriptive analysis of workaround origins and their effects on system design and data integrity. Our approach provides insights into user behaviour during misalignments, advancing research in workaround detection and user-centered design. Analysis of these workarounds can inform automated detection mechanisms and recommendation systems for stakeholders in software design and usage.
2 Background and Related Work
Gasser defines workarounds in computing as follows [9]: “working around means intentionally using computing in ways for which it was not designed or avoiding its use and relying on an alternative means of accomplishing work.” They emerge for various reasons, such as limitations in computing systems, challenges in existing workflows, or the need for quick solutions [2]. Users often improvise when encountering technology misfits or data misalignments to complete their regular tasks. In our context, a technology misfit is a mismatch between the technology in use and the practical needs or contingencies of everyday work. Data misalignment occurs when application processes or forms do not align with real-world data. Workarounds are often seen as examples of bricolage and/or improvisation: Bricolage “involves making do with available resources, occurring over short or long periods” [10]; Improvisation “refers to bricolage in a short timeframe” [13].
Some workarounds are executed and disappear once the obstacle is overcome; others become informally embedded into organizational routines [23]. Often, institutionalized workarounds become stepping stones for planned changes to processes and software [2]. Researchers have thus attempted to systematize and understand them. A workaround ontology was developed by Röder et al. [17], while Zainuddin et al. proposed a taxonomy [28].
The data generated during the operation of information systems has been extensively leveraged to detect workarounds, e.g., by mining logs using deep learning methods [22], or for process mining [4, 20]. A common approach is to track the sequence of actions and keystrokes users take when interacting with software systems. Workarounds can then be identified by comparing intended processes to actual user behaviour. Other studies focus on understanding the impact of workarounds, noting that users are often unaware of the potential consequences when they engage in these behaviours [8, 16]. Additionally, Schou et al. highlighted how unnoticed workarounds can have negative effects on organizational processes [18].
Detecting workarounds in workflows often relies on interviews and case studies. Soffer et al. investigate the motivations for workarounds in processes and find that they arise from misalignments between user goals and official processes or conflicts among the goals themselves [19]. Van der Waal et al. investigated how workarounds evolve over time and whether they are adopted by individuals or groups, emphasizing the need for continuous monitoring [21]. Davidson et al. explored the reasons behind workarounds, such as inadequacies in information systems, where users bypass formal processes simply to get their work done [6]. Wibisono studied how workarounds affect data quality, noting that users may interpret data in ways that introduce subjectivity and result in pseudo-data quality [24]. Davidson et al. examined the role of no-code/low-code [7] tools in enabling workarounds that deviate from established organizational processes [5]. Mubarkoot et al., taking the perspective of software compliance, highlighted that while end-user workarounds have been studied, there is a lack of research on workarounds for software engineers [14].
3 Studying Workarounds
We aim to analyze the workarounds users do when interacting with forms in the context of misalignment between data and form software interfaces. We want to characterize their impact on user effort and data quality, with respect to frequency, and systematic behaviour.
By designing experiments using our proposed experimental method, researchers can investigate workarounds by examining their types, frequency, and relationships with each other and various types of misalignments.
Another important aspect to explore is the association between form widgets and the workarounds users attempt with them. This helps characterize different workaround strategies based on the available widgets users interact with when navigating software forms or workflows.
Additionally, researchers can analyze recurring patterns and systematic approaches users adopt when implementing workarounds. Another possible investigation is whether the frequency of misalignments affects the effort users invest in finding workarounds.
3.0.1 Experimental instruments
When investigating workarounds, researchers must clarify three key factors before conducting experiments: the creation of forms used in the study, the input data, and the assignment of users. Another critical instrument in workaround studies is a logger integrated into forms and workflows to capture user interactions, including widget usage, timing information, keystrokes, and mouse clicks. These logs are essential for analyzing interactions with form widgets and application navigation during workarounds.
For controlled or pilot experiments, researchers can use synthetic forms, data, and scenarios. They are responsible for recruiting and assigning users while designing various types of data misalignment, where input data does not fit the provided forms and workflows. Each form dataset should include both easily enterable records (where data format and content match expectations) and misaligned records.
There are multiple form widget categorizations, such as “action”, “data”, and “static” [11] or “navigation”, “input”, “control” and “presentation” [12]. We propose a categorization based on a widget’s affordance for user workarounds, i.e., the degree of freedom it offers for data manipulation. Open: Widgets that allow users to freely modify input data, such as text boxes and text areas with minimal or no validation rules. Predetermined: Widgets containing preset data, requiring users to select from existing options, such as dropdown lists, combo boxes, radio buttons, checkboxes, and colour pickers. Controlled: Widgets that accept data within a constrained (but potentially large) range, such as sliders, number spinners, file uploads, and date pickers. When designing forms, researchers should balance the distribution of widgets from these categories to fit their controlled experiment.
Although the results of such experiments may not be fully generalizable, they provide valuable insights and build confidence in conducting further empirical studies, such as case studies or natural experiments. In a case study, researchers do not create the forms but advise the case study partner on software form and workflow design. There is no synthetic data, but researchers may provide guidance on user assignment. In a natural experiment, researchers do not influence form design or user assignment. Their role is limited to integrating the logger, which captures user interactions within the software forms and workflows.
These studies contribute to understanding workarounds and improving their detection, ultimately helping to identify missing requirements, validate software forms and workflows, and even automate pull request generation for missing features.
3.0.2 Experiment Process
In controlled experiments and case studies, each participant requires an individual session to enter their data into the system. Sessions consist of three parts: introduction, tasks, and exit survey. During the introduction, researchers present experimental tasks and provide documentation about the software forms or workflows, including screenshots of completed forms to familiarize participants with the interface and key features. Participants should remain unaware of the experiment’s focus on workarounds and receive instructions only to enter data through the forms and workflows. While loggers capture quantitative data, they may miss certain behaviours, necessitating the exit survey. After completing data entry tasks, participants must complete this survey about their experience. The survey can include questions about the participant’s occupation or field of study, self-assessment of difficulty entering misaligned data, self-reporting of workaround usage, and open-ended questions explaining employed workarounds or suggesting improvements to handle misalignments.
3.0.3 Metrics
In controlled experiments and case studies where misalignments and widget forms are known, researchers must label records that exposed a misalignment (i.e., for which a workaround is expected) according to the workaround type and quality. They also collect the widget involved and its widget category (open, predetermined, or controlled).
Workaround type reflects how users adapt to misalignments [28], while workaround quality assesses data integrity and semantic consistency between original and entered data [27]. Quality scores range from 1 to 4: A score of 1 is significant data loss, a score of 2 is partial data loss, a score of 3 is indirect data retention (entered elsewhere or reported differently while preserving intent), a score of 4 is prominent input of misaligned data, and no score when misaligned data completely skipped (no workaround).
From record save times captured in the logger, researchers can deduce workaround implementation time. Since participants have unique data entry paces, researchers must calculate a distribution-sensitive baseline for normal record save times for each participant. After testing the distribution using Shapiro-Wilk: For a normal distribution, they can use the mean; for left/right-skewed distributions, the 30th/70th percentile, respectively; for heavy-tailed distributions, the trimmed mean (excluding the lowest/highest 10%); and for other cases, they can use the median.
Another suggested metric is relative workaround effort which represents time to complete a workaround adjusted to baseline, computed as the ratio of workaround save time to participant’s baseline.
Also, systematicity score indicates how consistently participants apply the same workaround type for the same misalignment type, calculated as the proportion of instances where the most common workaround was used, averaged across misalignment types.
4 Pilot Experiment
We conducted a sample pilot experiment [26] to instantiate our proposed method for studying user workarounds in software applications by analyzing interactions with forms for entering misaligned data. We used convenience sampling, sending invitations to research groups and communities within our network. No prior knowledge of any specific software was required, and 15 participants completed the experiment. They were tasked with entering prepared datasets into instrumented software applications we developed. Across the 3 tasks, each participant encountered at least 4 misalignment types for a total 45 misaligned data.
This pilot experiment allowed us to assess both the feasibility of our experimental method and the concepts we proposed for investigating workarounds. While our findings are not generalizable due to the small sample size, they provided valuable insights.
We observed when users encountering misalignments resorted to workarounds, how these workarounds affected data quality, and how users prioritized their strategies based on the quality of the final data they could enter. We also gained experience with our widget categorization: Open widgets provided flexibility, enabling users to enter unstructured or misaligned data; Predetermined widgets were more prone to data loss due to their rigidity; and how Controlled widgets influenced workaround strategies and data quality. Additionally, we explored whether users developed systematic strategies after repeatedly encountering the same type of misalignment.
Our results indicate that applying this experimental method, along with our proposed research concepts, metrics, and recommendations on instruments and process, can yield valuable insights into workarounds. Conducting larger experiments with more participants, case studies, or natural experiments based on our experimental method would be a worthwhile next step.
4.1 Threats to Validity
Our workaround quality measurement, though based on data retention to reduce subjectivity, may not fully reflect true quality. The systematicity score assumes repeated use of the same workaround type indicates intentional behaviour, potentially overlooking factors like user background or problem interpretation. Participants might attempt more workarounds than usual due to the Hawthorne effect [1]. To mitigate this, we carefully worded experiment instructions to focus on data entry and software interaction without mentioning workarounds.
Despite task and misalignment randomization, earlier tasks may unintentionally improve performance in later ones. The tasks and application domains in our pilot experiment cannot fully represent real-world scenarios or capture the complexity of real-world misalignments, limiting result generalizability. While forms provide a suitable focus for this research stage given their pervasiveness in software, this scope may limit findings’ applicability to systems where workarounds differ.
Our participants’ engineering and computer science backgrounds may have influenced their approach despite being assigned non-technical tasks. For instance, computational thinking [25] might affect workaround types employed. Future experiments should include more diverse participants. Researchers adopting our experimental method should consider participant background diversity.
Finally, our interpretations of participants’ workaround choices may involve subjective judgment. We mitigated this by interpreting results and then cross-referencing participants’ exit survey responses to corroborate findings.
5 Summary and Discussion
We proposed an experimental method to study how users employ workarounds when software forms do not align with their data or goals. A pilot experiment validated its feasibility and research questions. Applying this method provides insights that support future research and help designers better address user needs. Future work could extend this approach to non-form-based systems and explore automated workaround detection to improve user-centred software.
Experiments based on this method can reveal patterns in how users handle form data misalignments. Researchers can determine whether users choose workarounds over abandonment, prioritize data integrity in their workaround strategies, or combine multiple workarounds. Another key aspect is whether users become more efficient after repeatedly encountering misalignments. Additionally, our widget categorization provides a framework for examining the relationship between workaround quality and widget affordances.
Workarounds highlight design flaws but also pose risks to data integrity. Software designers must proactively understand workarounds to identify unmet needs and mitigate risks. Survey results from these experiments can help researchers assess whether detecting workarounds uncovers design problems or confirms that misaligned designs push users toward shadow systems or complete bypass [9].